OpenStudy (anonymous):

Consider the area between the graphs x+3y=1 and x+9=y2. This area can be computed in two different ways using integrals First of all it can be computed as a sum of two integrals; Alternatively this area can be computed as a single integral

OpenStudy (amistre64):

and the question for this statement?

OpenStudy (amistre64):

we can integrate them both at the same time, or seperately and add the sums...

OpenStudy (anonymous):

I need both please!

OpenStudy (anonymous):

And I need a and b, as well as c

OpenStudy (amistre64):

the common line between them is the sum of there parts./2 right? gets an averge?

OpenStudy (anonymous):

I also need f(x) and g(x) separately

OpenStudy (amistre64):

f(x)=-x/3 +1/3; g(x) = sqrt(x+9)

OpenStudy (amistre64):

[S] f(x) dx - [S] g(x) dx [S] f(x) - g(x) dx

OpenStudy (anonymous):

The fist thing you have to do is to find the intersection of the two lines. Secondly, you have to find the intersection of the two lines with the x axis. As the third step, you have to integrate the frist integral from the x-coordinate of the first line until the x-cordinate of the intersection point. You have to do the same with the seconde line

OpenStudy (amistre64):

it doesnt provide an interval, and so assuming that it just wants the intersections is presumptuous; prolly true, but not accurate

OpenStudy (amistre64):

it could want the interval [0,root to the right]

OpenStudy (anonymous):

Im not sure how to do that. Sorry

OpenStudy (amistre64):

you get something that looks like this

OpenStudy (amistre64):

without a stated interval, you are prolly correct in assuming the left and right intercepts

OpenStudy (anonymous):

ahhh ok. give me one second to thing about it

OpenStudy (amistre64):

its got issues of being on the other side of the y axis; and youd have to split it into intervals from the common point to the left and right

OpenStudy (amistre64):

maybe +9 to get it all on the right side of the y axis?

OpenStudy (amistre64):

and my line parts to steep , but its just an artist interpretation lol

OpenStudy (amistre64):

kinda like this eh

OpenStudy (anonymous):

Anyway, you have to split the total area in two sides.Carry out integration from intersection of the parable and the x-axis until the intersection of the parable and the line. They you have to integrate from the intersection point to the intersection of the line with the x-axis. That is how I would do that my friend

OpenStudy (anonymous):

I got x=1 for c, but it keeps saying that it's wrong.

OpenStudy (anonymous):

which c?

OpenStudy (anonymous):

the end point on the right

OpenStudy (anonymous):

I dont know my friend. I got a little bit confused. Sorry :(